A sample of cisplatin is 65.02% platinum, 9.34% nitrogen, 2.02% hydrogen, and 23.63% chlorine. Determine the empirical formula.


1
Expert's answer
2022-04-27T14:53:02-0400

Solution:

Step 1. Determine atomic mass weighted values.

Pt: 65.02/195.08 = 0.3332

N: 9.34/14 = 0.6671

H: 2.02/1 = 2.02

O: 23.63/35.45 = 0.6665

Cisplatin, the common name for a platinum compound that is used to treat cancerous tumors, has the composition (mass percent) 65.02% Pt, 9.34% N, 2.02% H and 23.63% Cl. Calculate the empirical formula for cisplatin. Step 2. Determine the empirical formula by taking ratios to the smallest value. N:Pt = 0.6671/0.3332 = 2.01 H:Pr = 2.02/0.3332 = 5.98 H:N = 0.6665/0.3332 = 2.00 The formula is PtCl2N2H6